OpenStudy (anonymous):

Help! 2x/x+2 + x/x-2 - 1/x^2-4

OpenStudy (anonymous):

\[\frac{ 2x }{ x+2 }+\frac{ x }{ x-2 }-\frac{ 1 }{ x^2-4 }\]